OpenStudy (anonymous):

Solve 9(g-10)-4g=8g+27 with steps?

OpenStudy (anonymous):

distribute 9 over (g-10)

OpenStudy (kropot72):

Multiplying out: 9g - 90 - 4g = 8g + 27 Grouping like terms: 9g - 4g -8g = 27 + 90 -3g = 117 \[g=-\frac{117}{3}\]

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

You're welcome :)